OpenStudy (chris215):

Va is the orbital speed of a satellite at a distance x from the centre of the Earth. If the satellite is at a distance 2x from the centre of the Earth and with the speed of Vb, which of the following statements is correct?

OpenStudy (chris215):

va=vb^2

OpenStudy (anonymous):

well the answer is Vb = Va^2 Va^2 = Gm/r and Vb^2 = Gm/2r which leads you to get 2(Vb^2) = Va^2 which gets you Vb = Va^2

OpenStudy (chris215):

these are the only answer choices I have

OpenStudy (anonymous):

this is physics right?

OpenStudy (chris215):

yeah i already posted it in physics and no one has helped

OpenStudy (loser66):

Is it not that @Marcorie gave you the answer which is b) ??

OpenStudy (loser66):

He was downhill without apply the break, so that he had one more extra simplification and that confused you!! stop at \(2V_b^2=V_a^2\) , which is option b

OpenStudy (loser66):

*applying